OpenStudy (anonymous):

25. Solve the equation. 6 =x+2/3 (2 points)

OpenStudy (anonymous):

x=16/3

OpenStudy (anonymous):

x+2/3=6 x=6-2/3 x=18/3-2/3 x=16/3

OpenStudy (anonymous):

any other problems?